Write the balanced equation for the reaction of aqueous Pb(ClO3)2 with aqueous NaI. Include phases. chemical equation: What mass of precipitate will form if 1.50 L of highly concentrated Pb(ClO3)2 is mixed with 0.300 L 0.260 M NaI ? Assume the reaction goes to completion. mass of precipitate:
Answer #1

Balanced Equation:'

Pb(ClO3)2(aq) + 2 NaI(aq) PbI2(s) + 2 NaClO3(aq)

From the equation, we see that: moles of PbI2 = moles of NaI / 2

Moles of NaI = molarity * Volume = 0.260 * 0.30 = 0.078 moles

Moles of PbI2 = 0.078 / 2 = 0.039 moles

Mass of PbI2 = moles * molar mass = 0.039 * 461 = 17.98 g